OpenStudy (anonymous):

Mia has taken 3 more tennis lessons than Rita has taken. Mia has taken 8 tennis lessons. What is the total number of tennis lessons that Mia and Rita have taken? lessons

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

5? but it could be something else too :P

TheSmartOne (thesmartone):

So if Rita has taken x lessons then Mia has taken x+3 lessons

563blackghost (563blackghost):

Mia has taken 8 lessons and rita toke 3 less lessons so 8-3=?.

OpenStudy (anonymous):

Mia has taken more lessons then Rita @TheSmartOne

OpenStudy (anonymous):

8-3=5.. i just wanted to double cheek because i want sure :P

563blackghost (563blackghost):

What is the total number of tennis lessons that Mia AND Rita have taken? so i guess you have to add them up together.

OpenStudy (anonymous):

13 its how much altogether

TheSmartOne (thesmartone):

Because then Mia has taken 8 And Rita has taken x+3 =8 and so x=5

TheSmartOne (thesmartone):

And the question wants to know how much together.

TheSmartOne (thesmartone):

So you add \(\sf 8+5=?\)

OpenStudy (anonymous):

13

TheSmartOne (thesmartone):

So that is your answer :)